OpenStudy (anonymous):

can anyone help me slope intercept form 2x+3y=9

OpenStudy (anonymous):

Slope intercept form => y = mx + b where m is the slope b is the y-intercept Transform the above equation to look like that.

OpenStudy (anonymous):

3y = -2x _9 divide by y = (-2/3)x - 3 slope = -2/3 intercept is -3